### v1

# Dimensionality increase for error correction in the interaction between information space and the physical world

Abstract. The evolution of human intelligence led to the huge amount of data in the information space. Accessing and processing this data helps in finding solutions to applied problems based on finite-dimensional models. We argue, that formally, such a mathematical model can be embedded into a higher-dimensional model inside of which a desired solution will exist.
Keywords: information, data, errors, mathematical models, manifolds, processes, dimension
1. Introduction
There is a vast amount of literature on the computational and mathematical methods of learning (e.g. the monograph [11]), the interaction between a computer and a human, and theoretical aspects of information engineering related to signal processing or control theory. Geometric methods, especially those based on Riemannian geometry, have been extensively used in the study of networks, information transmission, and data analysis: see e.g. discussion and references in [2], [3].
In [15], it is stated that the three stages of development of human intelligence are based on
Stage 1: interaction between a person and the physical environment, (e.g. concluding that the fire is hot);
Stage 2: primitive interaction between a person and another human (e.g. repeating another person’s actions such as drinking water);
Stage 3: a higher level interaction among humans based on language, i.e. exchange of information that involves encoding and decoding processes.
The author of [15] draws parallels between these stages and machine learning. In particular, he compares Stage 1 to unsupervised learning.
In Stage 3, the encoded information exchanged among people would be primarily about the physical world, themselves (as part of this world), and related processes. By now, humans created a huge repository of data, in various formats. This information/data is typically supported by an object in a physical world (e.g. a printed book, a painting, an mp3 music file physically stored on a computer drive). Communication among people now involves new information that addresses (encodes) only information created by people. It is information about information (about abstract information, rather than about physical world). As examples, that includes techniques of solving equations, certain algorithms, or machine learning methods.
Predicting, analyzing and correcting errors (or malfunctions) is an essential part of data analysis and engineering. To give some specific examples, we mention the paper [9] that discusses the contributing factors to errors in machining and presents theoretical and quantitative analysis of the model that leads to improved accuracy. The word ”error”, in the most general sense, can be understood as an issue to be fixed or a problem to be addressed. In the abstract sense, it can be parameters or data that need to be modified. The paper [10] reviews the issues that occur in large networks, such as congestion, delays or packet loss, and possible approaches to gaining control over these problems. The paper [4] addresses the problem of low power in radio frequency energy harvesting systems. Mathematical representations or models of real life systems often serve as useful approximations. For instance, in [12], the model for an electrical vehicle system allows to compare and evaluate the system responses.
2. Main theorem and proof
In this section, we describe our mathematical model.
Let H be an infinite-dimensional Hilbert space (real or complex). If H is separable, then it is unique up to an isomorphism. Equip H with the structure of a Hilbert manifold ([5], [6]). For our analysis, let the physical world be a submanifold of a manifold W=H, and let the information space be a submanifold of the manifold E=H. We will model a person (an individual human being) by a manifold M, together with a probability measure, p, on M. In a one-parameter smooth family (M(t),p(t)) of such pairs (we can assume this to be a Bayesian process), where t is time, the choice of the probability measure p=p(t) is a ”randomizer”, which, informally speaking, reflects the unpredictability of the human choices. We assume the pair (M(t),p(t)) at every time t is represented as an embedded submanifold of W (i.e. the choice of p is incorporated in the embedding). As a simple example, in Fig. 1, a manifold (c,d) (the open interval $c<x<d$ in the x-axis), with a probability distribution p whose graph is shown in the picture, is embedded into the 2-dimensional plane as the red curve, the graph of p (rather than a line segment). This is for one fixed moment of time, say t=0.
Figure 1.
Generally, following the spirit of [2], [3], whenever possible, assume that everything relevant is already contained in the manifold (or parameter space) and it is unnecessary to deal with a choice of extra structure (e.g. a metric). For example, for the manifold $J=(c,d)$ (the line segment in Fig. 1) equipped with a choice of a metric $\rho$, a function on $J\times J$ that assigns to each pair $(x_{1},x_{2})$ the value of the distance between them, $\rho(x_{1},x_{2})$, we can replace the pair $(J,\rho)$ by the set of points in the xyz-space
$X=\{(x,y,z)|\ c<x<d;c<y<d;z=\rho(x,y)\}.$
Consider the semigroup G formed by mappings
$U\to V$
$V\to C\to U$
$V\to C_{1}\to...\to C_{k}\to U$,
where $U$, $V$ are (any) submanifolds of W, and $C$, $C_{1}$,…, $C_{k}$, for a positive integer k, are (arbitrary) submanifolds of E. These maps include, in particular, the maps that represent direct human-to-human interactions, or the human-to-human encoded communication of any form, or a person putting data into E or retrieving information from E.
Everything in the information space is a result of processes in someone’s neuronal networks. Joining the information space to the physical world, in some sense, we are looking at the collective insides of our brains. A car, a bicycle, sports (e.g. basketball), - all this originated in someone’s mind, mapped to E, as a set of images and other data, and then realized in W, via the semigroup G.
The term ”information space” that we chose for our submanifold of E is not the same as ”information manifold” used in information geometry (see e.g. [1]). In information geometry, an information manifold (or a statistical manifold) is a set of probability measures on a measure space, equipped with additional structure (e.g. Fisher information metric). It is possible, under appropriate assumptions, to make it a Hilbert manifold [8]. But this is not necessarily the manifold H in our context. Our H is an abstract infinite-dimensional Hilbert space with an infinite-dimensional manifold structure.
Consider the question of fixing a malfunction (or correcting an error) in W. This is a broad question that can include actions such as replacing a failing battery or giving a person a medication.
Theorem 1. Let $W$, $E$ be as defined above. Let $a<b$ be real numbers and let $I=[a,b]$. Let $n>1$ be the dimension of $S$, the submanifold of $W$ that models the physical world. Let $A$ be a submanifold of $S$. Assume the manifolds $S$, $A$ are compact, with or without boundary. Let $B$ be a finite-dimensional submanifold of $W$, diffeomorphic to $A$, such that $A\cap B=\emptyset$. Then there exist a submanifold $X$ of $W$ of dimension $n\leq\dim X\leq 2n+1$ and a smooth family of submanifolds $Y(t)$, $t\in I$, of $X$ such that $Y(a)=A$ and $Y(b)=B$.
Proof. We have: $\dim A\leq\dim S$, $\dim B=\dim A$. Since $S$ and $B$ are compact finite-dimensional submanifolds of $W$ and $\max\{\dim S,\dim B\}=n$, it follows that there exists a compact n-dimensional submanifold $K$ of $W$ that contains both $S$ and $B$. By the Whitney embedding theorem ([13], Theorem 1), there is a smooth embedding $f$ of $K$ into the $(2n+1)$-dimensional Euclidean space. Moreover, $f(A)$ and $f(B)$ are embedded submanifolds of ${\mathbb{R}}^{2n+1}$. Let’s identify this ${\mathbb{R}}^{2n+1}$ with a finite-dimensional linear subspace of $E$. We have: $f(K)$ is compact in ${\mathbb{R}}^{2n+1}$ (image of a compact set under a continuous map), $f(A)\subset f(K)$, $f(B)\subset f(K)$ and $f(A)$ is diffeomorphic to $f(B)$. By the isotopy version of the Whitney embedding theorem ([14] if $\dim A=n$, Theorem 6/deformation theorem in [13] if $\dim A<n$) there is a smooth family of submanifolds $\Lambda(t)\subset{\mathbb{R}}^{2n+1}\subset E$,  $a\leq t\leq b$, such that $\Lambda(a)=f(A)$ and $\Lambda(b)=f(B)$. Let $V$ be an open $(2n+1)$-dimensional ball that contains $f(K)$ and the compact manifold
$\{(t,x)|\ a\leq t\leq b;\ x\in\Lambda(t)\}.$
Next, we claim that there is a smooth map $F:V\to W$ such that $F\Bigr{|}_{f(K)}=f^{-1}$. This follows from the proof of Lemma 2.2/extension lemma in [7]. Lemma 2.2 can not be used as stated in [7], because $W$ is infinite-dimensional, and Lemma 2.2. is for maps with codomain ${\mathbb{R}}^{m}$. Note that we can not assume that there exists a finite-dimensional linear subspace of $W$ that contains $S$ or $K$ (this may not be true). However, the proof of lemma 2.2 [7] can be adjusted, to the case of codomain $W$, to conclude that the map $f^{-1}:f(K)\to K$, regarded as a function $f(K)\to W$, extends to $F:V\to W$. We observe: $F(f(A))=A$, $F(f(B))=B$. We set $X=F(V)$ and $Y(t)=F(\Lambda(t))$ for each $t\in I$. The desired conclusion follows. $\Box$
The sketch in Fig. 2 illustrates the idea of the proof.
Figure 2.
3. Discussion
In this section, the example in subsection 3.1 is directly relevant to Theorem 1. Subsections 3.2, 3.3, 3.4 are meant to illustrate at an intuitive level how solving a problem is tied with increasing the number of parameters in the mathematical model that approximates this system.
3.1.
Deconvolution is a mathematical technique that sharpens blurry images. Typically, there is a signal recorded with an optical instrument (e.g. a microscope). This recorded signal is a function $\psi$ (the blurry image), and $\psi=\alpha\ast\beta$, where $\beta$ is a point spread function. The objective is to recover the true signal $\alpha$ (the image with a good resolution and contrast). There are various software packages commonly used for practical purposes, e.g. Fiji/ImageJ.
Applied to this situation, Theorem 1 can be interpreted as follows. Let A be a blurry photo image (represented by the signal $\psi$). The theorem asserts that there must exist a process (which possibly involves doubling the number of variables, plus one more variable) which results in a more clear photo, B. In our setting, B is represented by $\alpha$. One possible way to get from A to B is to by implementing deconvolution. Using the Fourier transforms amounts to doubling the number of variables. One extra variable is time. This is consistent with the assertion of the theorem.
3.2.
In a very basic information processing example, a bit flip error in a specific transmission of a string of 0s and 1s can be diagnosed by sending multiple copies of the same string through the same channel. In the quantum information set-up, this trick no longer works, due to the no-cloning theorem, but there are other ways to detect the bit-flip errors via increasing the number of qubits. Without going into the general theory of error correction, let us simply note that in both these examples, the intent to correct the error led to increasing the size of the space of states.
3.3.
In Fig. 4, there is a tennis ball held by two rods (see also Fig. 3). The xy-plane is vertical. The photo on Fig. 5 is taken after the ball is released. Suppose we predicted that upon being released, the ball would move in the negative direction of the y-axis. If we interpret it as a prediction error, then one way to correct it would be to add one more parameter to Fig. 3: the direction of the gravitational force, and this would allow us to make the correct prediction for the direction of the motion of the ball. Another way to add parameters to the situation would be to place a mechanical arm that would hold the tennis ball and, at time t=0 would move it in the prescribed direction (in this case, in the direction of the negative y-axis).
Figure 3.
Figure 4.
Figure 5.
3.4.
In Fig. 6, there is a photo of a broken pencil. We can edit the image to ”repair” the pencil: Fig. 7. This amounts to rearranging the pixels in a 2-dimensional submanifold of E. Mirroring this process in W means putting the two pieces together. However, the actual pencil in W is still broken. The bonds at the atomic scale have not been repaired. The problem needs to be formulated with more parameters, to be solved in W.
Figure 6.
Figure 7.
4. Conclusions
The information space is distinct from the physical world. We model both by submanifolds of infinite-dimensional manifolds W and E, respectively. The information space contains the information created by processes in neural networks of human brains. Many processes can be modelled by mappings between subsets of W and subsets of E. For example, the idea of using digital twins to solve problems in the physical world fits perfectly in this setup. This direction of thinking is also consistent with world models in artifical intelligence.
Theorem 1 addresses the following situation. We consider submanifolds A and B of W, where A is in S, which we take to be the model of the physical world, or a relevant subset of the environment. The manifold S may contain, for instance, submanifolds that represent humans or other processes. The submanifold A represents a set of parameters that contains an ”error” (a localized data set that needs to be changed) and B the target set of parameters where this error is not present. As an informal comment, in practice, fixing one local error may change the values of other parameters elsewhere, but we assume that B is an acceptable outcome. We consider the Bayesian process in S, over the time interval I=[a,b], starting with A at t=a. Because it is governed by the laws included in the definition of S, the process will finish at a submanifold of S at time t=b. The theorem does not require that B is a subset of S. The assumption is that B is a submanifold of W. If A is allowed to evolve on its own, within the limitations set by the model and the present conditions, then the outcome B may not be reachable. The theorem asserts that it is possible to find a smooth path from A to B, and that is comes with increasing the number of parameters used to describe the physical world, S, from n parameters to (up to) 2n+1 parameters. Informally speaking, this maybe be regarded increasing the magnification/resolution. In other words, the theorem says that the error (or a malfunction) can be fixed by increasing the dimension of the model, but the theorem does not suggest how to choose these additional parameters (up to n+1 of those). The proof indicates that the extra parameters appear from mappings between subsets of W and subsets of E, i. e. elements of the semigroup G, discussed earlier. However, the proof does not specify which specific elements of G give these extra parameters. In this sense, the proof is not constructive and the theorem is an existence statement.
In Section 3, we present several examples that echo the thinking behind this theorem. The discussion of deconvolution in image processing demonstrates the statement of the theorem most precisely among these.
The theorem is meant to be a general statement/principle, applicable to mathematical models of image processing or signal processing in biomedical systems, or other systems or networks.